# ST-USleepNet: A Spatial-Temporal Coupling Prominence Network for Multi-Channel Sleep Staging

### v1

@@ 4.2 U-shaped Sleep Network
The spatial-temporal graph $G_{s}$ constructed from the raw signals is then input into the U-shaped sleep network (USleepNet). There are two interwoven networks, the temporal prominence network for the temporal stream and the spatial prominence network for the spatial stream in USleepNet.
@@ 4.2.1 Temporal Prominence Network
We propose the temporal prominence network for the temporal stream. This network consists of three logical submodules: the temporal encoder block $T_{en}$, the temporal decoder block $T_{de}$, and a segment classifier $T_{c}$. $T_{en}$ takes the feature matrix $X$ of the graph $G_{s}$ as input, extracts temporal features at different scales, and outputs a compressed feature representation. Then, $T_{de}$ continues to learn from the compressed feature representation provided by $T_{en}$ to obtain a point-wise label for each point. Last, $T_{c}$ integrates the point-wise labels into a segment-wise label.
Temporal Encoder Block
The temporal encoder block $T_{en}$ consists of two 2D convolution layers. The number of channels in the second convolution layer is greater than that in the first layer since deeper networks tend to capture more complex features. Each convolution uses the same padding, followed by batch normalization and dropout. The kernel size remains consistent within the two convolution operations of each $T_{en}$, while it may vary across different ${T}{en}$. After these two convolution operations, max pooling is performed to reduce the feature dimensions.
Temporal Decoder Block
The temporal decoder block $T_{de}$ consists of two transposed-2D convolution blocks. Each $T_{de}$ corresponds one-to-one with a $T_{en}$, using inverse operations to gradually recover the features. First, the compressed feature representation is unpooled. Next, to retain more detail and merge features of different scales, $T_{de}$ concatenates the unpooled vector with the output of the corresponding $T_{en}$. Next, the concatenated vector undergoes two convolution operations sequentially. $T_{de}$ shares some similarities with $T_{en}$, such as the kernel size remaining consistent within the two convolution operations of each $T_{de}$ but may vary across different $T_{de}$; and each CNN layer is followed by batch normalization and dropout layers.
Segment Classifier
The segment classifier $T_{c}$ is a module to get the segment-wise label, primarily composed of a fully connected layer. The input to ${T}_{c}$ is an embedding $H^{ch\times nC\times f_{s}t}$, where $ch$ is the channel number. First, $H$ is reshaped into ${\hat{H}}^{ch\times C\times nf_{s}t}$. Next, a softmax function is used to obtain the classification for each time point of each physiological signal channel, considered as point-wise labels. Then, mean pooling is used to integrate the channel $h$ into one, and a fully connected layer is used to obtain the final label.
@@ 4.2.2 Spatial Prominence Network
The input to $S_{en}$ consists of the adjacency matrix $A^{v\times v}$ and the feature matrix $X^{v\times h}$ of the spatial-temporal graph $G_{s}$, where $v$ represents the number of nodes and $h$ represents the feature dimensions. The first graph convolution layer maps $X$ to a higher-dimensional space, and the second graph convolution layer maps $X$ back to its original dimensions. The two consecutive graph convolutions can be represented as follows:
$\mathbf{X}^{(2)}=Drop(\sigma(\mathbf{A}Drop(\sigma(\mathbf{A}\mathbf{X}\mathbf{W}_{1}^{h\times\hat{h}}))\mathbf{W}_{2}^{\hat{h}\times h})$ (6)
where $Drop$ means the dropout layder, and $\sigma$ is ELU, which is defined as:
$\sigma(x)=\text{ELU}(x)=\begin{cases}x&\text{if }x>0\\ \alpha(e^{x}-1)&\text{if }x\leq 0\end{cases}$ (7)
Then, a 2D graph pooling layer (2D gPool) is used to perform down-sampling on the graph data. This reduces the size of the graph while retaining information on important nodes. In 2D gPool, we rank the nodes based on importance and select k to generate a subgraph. The computation of gPool can be represented as follows: First, each node’s feature vector $X_{i}$ is projected using a trainable projection vector $P$ to calculate a scalar projection value as an importance score $ic$:
$ic^{v}=\frac{X^{v\times h}p^{h}}{|p^{h}|}$ (8)
@@
$\text{idx}=\text{rank}(ic,k)$ (9)
$\tilde{ic}^{k}=sigmoid(ic(\text{idx})^{k})$ (10)
@@
$\tilde{X}^{v\times k}=X^{v\times h}(\text{idx},:)$ (11)
$\tilde{A}^{k\times k}=A^{v\times v}(\text{idx},\text{idx})A^{v\times v}(\text{idx},\text{idx})$ (12)
$\tilde{X}^{k\times h}=\tilde{X}^{v\times h}\odot(\tilde{ic}\mathbf{1}_{h}^{T})$ (13)
@@
Spatial Decoder Block The Spatial Decoder Block $S_{de}$ is used to recover the compressed subgraph obtained by $S_{en}$ into a feature supergraph. It consists of four network layers: a convolution layer, a 2D graph unpooling layer (2D gUnpool), and two graph convolution layers.
$S_{de}$ takes the adjacency matrix $A$ and the feature matrix $X$ of a spatial-temporal graph $G_{s}$ as input. First, a CNN layer is used to extract local information from the feature matrix $X$ and reduce the number of channels in $X$. Next, the data is upsampled using a 2D gUnpool layer. The 2D gUnpool layer corresponds to a 2D gPool layer and uses the node selection information recorded during the downsampling process of the 2D gPool layer. It places the existing nodes and their feature matrices back into their original positions in the graph, while the features of other nodes are set to zero. To avoid significant information loss, a skip connection is used to add the adjacency matrix of the upsampled graph to the corresponding adjacency matrix before downsampling in $S_{de}$. Finally, similar to $S_{en}$, two consecutive graph convolution layers are used: the first maps the feature matrix to a hidden high-dimensional space, and the second restores it to its original dimensions. Through these four layers, the $S_{de}$ block ultimately recovers the compressed feature subgraph into a feature supergraph.
@@ 4.2.3 Spatial-Temporal Block Fusion
The USleepNet is composed of an alternating combination of the temporal prominence network and the spatial prominence network. USleepNet includes $l$ layers of spatial-temporal encoders ${ST}_{en}$ and $l$ layers of spatial-temporal decoders ${ST}_{de}$, with a segment classifier at the end. Specifically, ${ST}_{en}$ consists of a $T_{en}$ followed by an $S_{en}$, and ${ST}_{de}$ consists of an $S_{de}$ followed by a $T_{de}$. In the encoder phase, processing with $T_{en}$ first is essential because PSG data usually has high dimensions and contains a lot of noise. Convolution and pooling operations can reduce the dimensions while removing noise, and extracting more representative features. These lower-dimensional features can reduce the computational complexity of subsequent graph processing.
@@ 5.1 Datasets
We evaluate the ST-USleepNet on three public datasets. In this part, We briefly describe each dataset and refer to the original papers for details.
ISRUC-S1 [25] contains PSG data from 100 patients (55 male and 45 female), with a total of 87,187 sleep segments, each lasting 30 seconds. The PSG data includes 10 channels of EEG, EOG, and EMG. ISRUC-S3 [25] contains PSG data from 10 patients (9 male and 1 female), with a total of 8,589 sleep segments. The PSG data includes 10 channels of EEG, EOG, and EMG. MASS-SS3 [26] contains PSG data from 62 healthy people (28 male and 34 female), with a total of 59,304 sleep segments. The PSG data includes 25 channels of EEG, EOG, EMG, and ECG.
@@ 5.2 Experiment Settings
To comprehensively evaluate the performance of ST-USleepNet, we select 9 models as baselines. These include classic works in sleep staging [27, 2, 4]; sleep networks based on U-shaped structures [3, 6]; spatial-temporal sleep networks [14, 24]; and the latest general multivariate time series model [28].
In the experiments, we divide the dataset into train, validation, and test sets, comprising 80%, 10%, and 10% of the total data, respectively. We select the best-performing model on the validation set and then test it on the test set. The evaluation metrics used are accuracy and F1 score.
@@ 5.3 Experiment Results
We compare ST-USleepNet with baseline models on three public datasets. The results are shown in Figure 1, where the best results for each metric are highlighted in bold and the second best is underlined. Our model achieves state-of-the-art performance across all metrics and demonstrates significant improvements compared to existing methods. On the three datasets, ST-USleepNet shows improvements over existing methods in accuracy by 1.7%, 4.2%, and 0.9%, and in F1 score by 1.8%, 6%, and 3.3%, respectively. This underscores the importance of extracting prominent sleep features and capturing spatial-temporal coupling patterns for sleep staging.
@@
Dataset ISRUC-S1 ISRUC-S3 MASS-SS3
Model Acc F1 Acc F1 Acc F1
RF 66.4 60.8 70.2 68.8 69.1 65.4
SVM 67.3 62.8 71.8 60.2 69.3 66.7
DeepSleepNet 73.0 69.1 75.7 73.2 84.8 79.8
TinySleepNet 75.3 73.9 76.1 74.5 82.4 77.8
U-time 71.8 67.1 75.5 71.9 85.0 79.3
GraphSleepNet 77.3 74.7 80.6 78.2 87.2 85.9
SalientSleepNet 80.6 80.0 80.3 78.4 84.9 82.8
BSTT 77.7 76.1 81.9 80.3 88.7 86.0
FC-STGNN 76.4 73.4 78.3 76.8 86.2 85.1
Ours 82.3 81.8 86.1 86.3 89.6 89.3
@@ 5.4 Ablation Study
To verify the effectiveness of the modules designed in ST-USleepNet, we design three variant models by removing the corresponding modules: 1) ST-USleepNet without temporal prominence network (-T). 2) ST-USleepNet without spatial prominence network (-S). 3) ST-USleepNet without spatial-temporal graph construction (-ST). In the ’-ST’ variant mode, we construct a spatial-only graph using the raw EEG signals in the whole segment based on the cosine similarity.
Figure 3: Results of ablation study on MASS-SS3 dataset.
Figure 3 shows the comparison results of the ablation experiments between the variant models and the original model on the MASS-SS3 dataset. It can be seen that removing either the temporal prominence network or the spatial prominence network adversely affects the model’s performance, and the impact is similar. This indicates that extracting characteristic sleep waveforms and salient spatial brain networks are both important. Additionally, removing the temporal-spatial graph construction module also affects model performance, suggesting that we should also focus on the spatial-temporal coupling patterns contained throughout the entire sleep signal, rather than just temporal or spatial features. Additionally, we explore the impact of important hyperparameters on the model’s performance. Detailed results are provided in the appendix.
@@ 5.5 Model Visualization
In the ST-USleepNet, each module can extract one kind of sleep feature: The temporal prominence Network can simultaneously capture characteristic sleep waveforms from multiple channels; The spatial prominence Network can capture the salient spatial Brain networks; The spatial-temporal graph construction module could capture the spatial-temporal coupling patterns. We show the visualizations for these three kinds of sleep features in this section.
Figure 4: Visualization results of extracted features.
@@ 5.5.1 Characteristic Sleep Waveforms
We visualize the model’s point-wise label output for each channel, demonstrating that the ST-USleepNet can simultaneously capture the characteristic sleep waveforms of various stages in multi-channel sleep data. As shown in Figure 4 (a), examples of characteristic sleep waveforms captured by the model for the five sleep stages are displayed. Each stage shows the results for two channels: for the W, N1, N2, and N3 periods, the F3 and C3 EEG channels are selected; for the REM period, the F3 EEG channel and the LOC EOG channel are chosen. It is evident that the model can simultaneously capture multi-channel wake period $\alpha$ waves; N1 period vertex waves, N2 period spindles, N3 period $\delta$ waves, and during the REM period, the EEG’s sawtooth waves and EOG’s rapid eye movement. In summary, our model effectively extracts characteristic sleep waveforms from multi-channel PSG data simultaneously.
5.5.2 Salient Spatial Brain Networks
We average the spatial-temporal graphs obtained from USleepNet into a single spatial graph and extract the salient brain networks according to the node of subgraphs output by the bottom block. The visualization results are in Figure 4 (b). In the figure, the left picture displays the complete brain network and several other nodes like EOG, while the right pictures show the extracted salient brain networks for each sleep stage. It is evident that the salient brain networks vary significantly across different sleep stages. More, specifically, during the wake stage, all brain regions are relatively active; in the N1 stage, as sleep onset begins, brain activity intensifies and eye movement reduces compared to the Wake stage; in the light sleep N2 stage, eye movement activity weakens further; in the N3 stage, deep sleep is entered, and brain network activity decreases; and during the REM stage, brain network activity is weaker, but eye movement is notably active. Thus, the ST-USleepNet can extract salient brain networks and possesses a certain degree of interpretability.

### v2

@@ 3.2 U-shaped Sleep Network
The spatial-temporal graph $G_{s}$ constructed from the raw signals is then input into the U-shaped sleep network (USleepNet). There are two interwoven networks, the temporal prominence network for the temporal stream and the spatial prominence network for the spatial stream.
@@ 3.2.1 Temporal Prominence Network
This network consists of three logical submodules: the temporal encoder block $T_{en}$, the temporal decoder block $T_{de}$, and a segment classifier $T_{c}$. $T_{en}$ extracts multi-scale temporal features from the feature matrix $X$ of the graph $G_{s}$, outputting a compressed representation, which $T_{de}$ further processes. Last, $T_{c}$ generates a segment-wise label.
Temporal Encoder Block $T_{en}$ consists of two 1D convolution layers. The number of channels in the second convolution layer is greater than that in the first layer, since deeper networks tend to capture more complex features. Each convolution uses the same padding, followed by batch normalization. The kernel size remains consistent within the two convolution operations of each $T_{en}$, while it may vary across different $T_{en}$. After these two convolution operations, max pooling is performed to reduce the feature dimensions.
Temporal Decoder Block $T_{de}$ consists of two transposed-1D convolution layers. Each $T_{de}$ corresponds to a $T_{en}$, performing inverse operations to recover features. It unpools the compressed representation, concatenates it with the corresponding $T_{en}$ output, and applies two sequential convolutions. Kernel sizes are consistent within each $T_{de}$ but may vary across blocks, with batch normalization applied after each convolution.
Segment Classifier $T_{c}$ generates segment-wise labels. It takes embedding $H^{ch\times nC\times f_{s}t}$ as input, reshapes it to ${\hat{H}}^{ch\times C\times nf_{s}t}$, and applies a softmax function to produce point-wise labels for each time step and channel. Mean pooling combines the channels into one, and a fully connected layer outputs the segment label.
@@ 3.2.2 Spatial Prominence Network
The input to $S_{en}$ consists of the adjacency matrix $A^{v\times v}$ and the feature matrix $X^{v\times h}$, where $v$ represents the number of nodes and $h$ represents the feature dimensions. The first graph convolution layer maps $X$ to a higher-dimensional space, and the second graph convolution layer maps $X$ back to its original dimensions. The two consecutive graph convolutions can be represented as follows:
$\mathbf{X}^{(2)}=Drop(\sigma(\mathbf{A}Drop(\sigma(\mathbf{A}\mathbf{X}\mathbf{W}_{1}^{h\times\hat{d}}))\mathbf{W}_{2}^{\hat{d}\times h})\text{,}$ (5)
where $Drop$ means the dropout layer, $W_{i}$ is a learnable hyperparameter, $d$ is the higher dimension than $h$, and $\sigma$ is ELU, which is defined as:
$\sigma(x)=\text{ELU}(x)=\begin{cases}x&\text{if }x>0\text{,}\\ \alpha(e^{x}-1)&\text{if }x\leq 0\text{.}\end{cases}$ (6)
Then, a graph pooling layer (gPool) is used to perform downsampling on the graph data. This reduces the size of the graph while retaining information on important nodes. In gPool, we rank the nodes based on importance and select k to generate a subgraph. The computation of gPool can be represented as follows: First, each node’s feature vector $X_{i}$ is projected using a trainable projection vector $P$ to calculate a scalar projection value as an importance score $ic$:
$ic^{v}=\frac{X^{v\times h}p^{h}}{|p^{h}|}\text{,}$ (7)
@@
$\text{idx}=\text{rank}(ic,k)\text{,}$ (8)
$\tilde{ic}^{k}=sigmoid(ic(\text{idx})^{k})\text{,}$ (9)
@@
$\tilde{X}^{k\times h}=X^{v\times h}(\text{idx},:)\text{,}$ (10)
$\tilde{A}^{k\times k}=A^{v\times v}(\text{idx},\text{idx})A^{v\times v}(\text{idx},\text{idx})\text{,}$ (11)
$\tilde{X}^{k\times h}=\tilde{X}^{k\times h}\odot(\tilde{ic}\mathbf{1}_{h}^{T})\text{,}$ (12)
@@
Spatial Decoder Block $S_{de}$ is used to recover the compressed subgraph obtained by $S_{en}$ into a feature supergraph. It consists of four network layers: a convolution layer, a graph unpooling layer (gUnpool), and two graph convolution layers.
$S_{de}$ takes the adjacency matrix $A$ and the feature matrix $X$ as input. First, a CNN layer is used to extract local information from the feature matrix $X$ and reduce the number of channels in $X$. Next, the data is upsampled using a gUnpool layer. The gUnpool layer corresponds to a gPool layer and uses the node selection information recorded during the downsampling process of the gPool layer. It places the existing nodes and their feature matrices back into their original positions in the graph, while the features of other nodes are set to zero. To avoid significant information loss, a skip connection is used to add the adjacency matrix of the upsampled graph to the corresponding adjacency matrix before downsampling in $S_{de}$. Finally, similar to $S_{en}$, two consecutive graph convolution layers are used: the first maps the feature matrix to a hidden high-dimensional space, and the second restores it to its original dimensions. Through these four layers, the $S_{de}$ block ultimately recovers the compressed feature subgraph into a feature supergraph.
@@ 3.2.3 Spatial-Temporal Block Fusion
USleepNet comprises an alternating combination of the temporal prominence network and the spatial prominence network. It includes $l$ layers of spatial-temporal encoders (${ST}{en}$), $l$ layers of spatial-temporal decoders (${ST}{de}$), a bottom block (${ST}{b}$), and a segment classifier at the end. Specifically, each ${ST}{en}$ consists of a temporal encoder ($T_{en}$) followed by a spatial encoder ($S_{en}$), while each ${ST}{de}$ consists of a spatial decoder ($S{de}$) followed by a temporal decoder ($T_{de}$). The bottom block (${ST}{b}$) is similar to ${ST}{en}$ but does not include any downsampling. In the encoder phase, it is essential to process with $T_{en}$ first, as PSG data typically has high dimensions and contains significant noise. Convolution and pooling operations not only reduce dimensions but also remove noise, thereby extracting more representative features. These lower-dimensional features significantly reduce the computational complexity of subsequent graph processing.
@@ 4.1 Datasets
We evaluate ST-USleepNet on three public datasets. ISRUC-S1 [34] contains PSG data from 100 patients (55 male and 45 female), with a total of 87,187 sleep segments. The PSG data includes 10 channels of EEG, EOG, and EMG. ISRUC-S3 [34] contains PSG data from 10 patients (9 male and 1 female), with a total of 8,589 sleep segments. The PSG data includes 10 channels of EEG, EOG, and EMG. MASS-SS3 [35] contains PSG data from 62 healthy people (28 male and 34 female), with a total of 59,304 sleep segments. The PSG data includes 25 channels of EEG, EOG, EMG, and ECG. For each dataset, the PSG data follows the AASM standards [6] and is classified into five categories: Wake, N1, N2, N3, and REM.
@@ 4.2 Experiment Settings
In this study, we use a cross-subject dataset division and perform five-fold cross-validation, ensuring that each data point is used in both the training and testing phases. The evaluation metrics used are accuracy (Acc) and F1 score.
@@
4.3 Baseline Methods
To comprehensively evaluate the performance of ST-USleepNet, we select 9 models as baselines. These include traditional Random Forest (RF) and Support Vector Machine (SVM) methods [36]; classic CNN and RNN-based methods DeepSleepNet [8] and TinySleepNet [10]; sleep networks based on U-shaped structures U-time [9] and SalientSleepNet [11]; spatial-temporal sleep networks GraphSleepNet [22] and BSTT [33]; and the latest general multivariate time series model FC-STGCN [37].
@@ 4.4 Experiment Results
We compare ST-USleepNet with baseline models in three public datasets. The results are shown in Figure 1, where the best results for each metric are highlighted in bold. Our model achieves state-of-the-art performance across all metrics. This underscores the importance of extracting prominent sleep features and capturing spatial-temporal coupling patterns for sleep staging.
@@
Model ISRUC-S1 ISRUC-S3 MASS-SS3
Acc F1 Acc F1 Acc F1
RF 73.81 70.43 73.99 70.20 79.35 77.09
SVM 75.05 72.79 73.56 71.88 81.65 78.23
DeepSleepNet 75.07 69.15 76.70 74.08 80.44 77.02
TinySleepNet 76.92 73.9 76.18 74.20 83.30 77.40
U-time 77.58 76.66 77.70 76.85 80.31 78.36
SalientSleepNet 78.83 78.16 79.34 78.54 85.30 82.35
GraphSleepNet 81.27 79.74 81.82 79.95 88.75 84.42
BSTT 81.96 80.30 82.18 81.54 89.50 85.00
FC-STGNN 77.34 76.77 80.87 78.79 86.23 85.19
ST-USleepNet 82.51 81.82 82.93 82.42 89.64 87.99
@@ 4.5 Ablation Study
To verify the effectiveness of the modules designed in ST-USleepNet, we design three variant models by removing the corresponding modules: 1) ST-USleepNet without temporal prominence network (-T). 2) ST-USleepNet without spatial prominence network (-S). 3) ST-USleepNet without spatial-temporal graph construction (-ST). In the ‘-ST’ variant model, we construct a spatial-only graph using the raw EEG signals in the whole segment based on the cosine similarity.
Table 2: Results of ablation study.
Ablation ISRUC-S1 ISRUC-S3 MASS-SS3
Acc F1 Acc F1 Acc F1
ST-USleepNet 82.51 81.82 82.93 82.42 89.64 87.99
-T 78.54 77.29 78.63 77.84 87.20 85.77
-4.81% -5.54% -5.19% -5.56% -2.72% -2.52%
-S 78.69 78.04 79.19 77.72 87.94 86.69
-4.63% -4.62% -4.51% -5.70% -1.90% -1.48%
-ST 78.35 77.79 78.52 78.08 87.49 86.07
-5.04% -4.93% -5.32% -5.27% -2.40% -2.18%
Table 2 shows the comparison results of the ablation experiments between the variant models and the original model on the three datasets. It can be seen that removing either the temporal prominence network or the spatial prominence network adversely affects the model’s performance, and the impact is similar. This indicates that extracting characteristic sleep waveforms and salient spatial brain networks is important. Additionally, removing the temporal-spatial graph construction module also affects the performance, suggesting that we should also focus on the spatial-temporal coupling patterns contained throughout the entire sleep signal, rather than just temporal or spatial features.
4.6 Hyperparameter Analysis
We explore the impact of various hyperparameters on the model’s performance and identify two with the most significant influence: convolutional kernel/hidden layer arrangement strategies and network depth.
First, ST-USleepNet implements multi-scale strategy in the temporal prominence network by dynamically varying convolutional kernel sizes and in the spatial prominence network by using different hidden layer dimensions, allowing the network to capture features at multiple scales effectively. To validate that the model adopts the optimal strategy, we experimented with the following combinations: 1)increasing then decreasing (our model’s strategy); 2)consistently using small kernels and hidden layers (small); 3)consistently using large kernels (large); 4)decreasing then increasing (reverse). We found that using decreasing convolutional kernels and hidden layer dimensions in the encoder, and increasing them in the decoder, achieves the best performance, as shown in Table 3.
Table 3: Comparison Across Different Convolution Kernel / Hidden Layer Arrangement Strategies.
Strategy ISRUC-S1 ISRUC-S3 MASS-SS3
Acc F1 Acc F1 Acc F1
Small 79.76 78.94 79.64 78.29 87.86 85.34
Large 80.16 79.43 80.96 79.46 88.47 86.56
Reverse 78.20 77.65 79.08 77.44 87.29 85.79
ST-USleepNet 82.51 81.82 82.93 82.42 89.64 87.99
Second, we investigate the impact of depth in the U-shaped network on the performance and determine that a four-layer network is the optimal structure, as shown in Table 4.
Table 4: Comparison of different network depths.
Depth ISRUC-S1 ISRUC-S3 MASS-SS3
Acc F1 Acc F1 Acc F1
2 76.41 75.28 76.26 75.56 86.17 85.57
3 78.97 77.49 79.14 78.88 87.78 87.16
4 82.51 81.82 82.93 82.42 89.64 87.99
5 82.60 81.75 82.87 82.43 89.22 87.88
@@ 4.7 Model Visualization
In ST-USleepNet, each module can extract one kind of sleep feature: the temporal prominence network can capture characteristic sleep waveforms from multiple channels simultaneously, revealing temporal patterns across different sleep stages; the spatial prominence network can capture the salient spatial brain networks, highlighting key brain regions involved in sleep; and the spatial-temporal graph construction module can capture the spatial-temporal coupling patterns, revealing how brain regions interact over time and space. In this section, we show the visualizations for these three types of sleep features, which not only provide insights into the underlying data patterns in different domains but also offer a level of interpretability for the predictive model.
Figure 4: Visualizations of sleep features that illustrates three types of sleep features across temporal, spatial, and spatial-temporal domains, highlighting the data patterns extracted by the model. This provides a degree of interpretability for the prediction.
@@ 4.7.1 Characteristic Sleep Waveforms
We visualize the model’s point-wise labels for each channel, showcasing that ST-USleepNet effectively captures characteristic sleep waveforms across multiple channels for various sleep stages. As shown in Figure 4 (a), the model identifies typical waveforms for the five sleep stages, with results displayed for two channels per stage. Notably, it captures $\alpha$ waves during the wake period, vertex waves in N1, spindles in N2, $\delta$ waves in N3, and during REM, EEG sawtooth waves and EOG rapid eye movements.
4.7.2 Salient Spatial Brain Network
We combine the spatial-temporal graphs generated by the model into a single spatial graph. From the bottom block’s output, we select channels with the largest sums of degrees across different time steps within the same channel. Using this information, we extract corresponding subgraphs from the concatenated spatial graph to represent the salient spatial brain network for each sleep stage. The visualization results are shown in Figure 4 (b). In the figure, the left panel shows the complete brain network, while the right panels display the salient brain networks for each sleep stage. These networks vary significantly: during Wake, brain regions are highly active; in N1, brain activity increases as eye movement decreases; in N2, eye movement further weakens; in N3, brain activity decreases significantly; and during REM, brain activity remains low while eye movement becomes strong.